OpenStudy (anonymous):

Find the determinant [1 3 1 ] 2 1 -5 3 -1 -4

OpenStudy (anonymous):

I got -35. is correct?

OpenStudy (anonymous):

yes that is correct :)

OpenStudy (anonymous):

Yea I think ur CORRECT

OpenStudy (anonymous):

Woo! Thanks!

OpenStudy (anonymous):

(1)(1)(-4)+(3)(3)(-5)+(1)(-1)(2)=-36 -3-5+24=16-------------->+16=20 |A|=20

OpenStudy (bahrom7893):

-35!

OpenStudy (anonymous):

@bnut you added up your first line wrong: (1)(1)(-4)+(3)(3)(-5)+(1)(-1)(2) = -51 -51+16 = -35

OpenStudy (anonymous):

oh shoot i did hahah what a boner